Suppose a producer faces the following production function: K^0,5.K^0,5 given that L is the unit of labour and K is the unit of capital. Suppose K is 16, labour wage rate is R10 and rental rate of capital is R50, derive the TC and MC functions


1
Expert's answer
2021-05-20T19:05:51-0400

Production function is given as:": Q=K^{0.5}L^{0.5}"

The amount of capital is fixed and equal to 16. So, the production function will look like:

"Q=16^{0.5}L^{0.5}"

"Q=4L^{0.5}"

"(\frac{Q}{4})^{2}=L"

"L=\frac{Q^2}{16}"

Total cost (TC) function of the firm is equal to:

"TC=w\u00d7L+r\u00d7K"

where w is wage rate, L is amount of labor, r is rental rate and K is amount of capital.

"TC=10L+16\u00d750"

"TC=10L+800"

Putting value of L in this equation:

"TC=\frac{10Q^2}{16}+800"

Marginal cost function is equal to:

"MC=\frac{\delta TC}{\delta Q}"

"MC=\frac{ 20Q}{16}"

"MC= \frac{5Q}{4}"